A quantity of Neon has a volume of 86.0 L at a pressure of 86.88 kPa and a temperature of 86.0˚C. What is the volume of the gas at STP, (101.3 kPa, 0.00˚C)?


1
Expert's answer
2021-04-07T11:28:16-0400

P1V1/T1 = P2V2/T2

P1 = 86.88

V1 = 86.0

T1 = 86.0 + 273 = 359K

P2 = 101.3

T2 = 273K

= 86.88×86/359 = 101.3 X/273

= 20.81248 = 0.371062X

X = 20.81248/0.371062

= 50.089

= 50.089 L